OpenStudy (vortish):

Identify the GCF and then factor the expression 5x^5+10x^4 -15c^3+10x^2

OpenStudy (vortish):

i think that the GCF is x^3

OpenStudy (vortish):

the 15c should be 15x

OpenStudy (mathstudent55):

Is that supposed to be -15x^3 or -15c^3? First look at the coeffeicients (the numbers): 5, 10, -15, 10 What is the common factor between them? Now look at the variable part: x^5, x^4, x^3, x^2 What factor do they have in common?

OpenStudy (mathstudent55):

Ok, I figured that.

OpenStudy (anonymous):

i would say the GCF is 5... @mathstudent55 what do u think?? is it 5 right?? cuz 5 is the largest thing that is in EVERY single component of the equation..

OpenStudy (vortish):

five is the common factor in all and x for the varible

OpenStudy (anonymous):

oh then nvm sorry just read the above thing... so GCF should be 5x^2 right??

OpenStudy (mathstudent55):

The GCF is made up of a number part and a varaible part. 5 is the number part. What is the variable part? How many x's do all terms have?

OpenStudy (anonymous):

yes, just follow what @mathstudent55 is saying ;) its all correct.. like his/her instructions

OpenStudy (mathstudent55):

All terms have at least x^2, so GCF is 5x^2

OpenStudy (vortish):

so we have the gcf now I have to factor the expression

OpenStudy (anonymous):

exactly :)

OpenStudy (vortish):

need help with that

OpenStudy (anonymous):

so based off of this, you can first factor out the GCF... do u understand what that means??

OpenStudy (vortish):

so we can 5x^2(all the rest of the expression)

OpenStudy (anonymous):

yup :) after that, what do you get?

OpenStudy (vortish):

heat do you have skype '

OpenStudy (mathstudent55):

5x^5+10x^4 -15c^3+10x^2 = 5x^2(_______ What is 5x^5 divided by 5x^2?

OpenStudy (anonymous):

no i don't.. sorry!!! all i have is OS and email lol

OpenStudy (vortish):

i c

OpenStudy (vortish):

5x^3

OpenStudy (mathstudent55):

Answer: x^3 5x^5+10x^4 -15c^3+10x^2 = 5x^2(x^3 ________ What is 10x^4 divided by 5x^2?

OpenStudy (anonymous):

@vortish, your GCF is 5x^2... and u divide, so like @mathstudent55 said, ur first part will be x^3 :)

OpenStudy (vortish):

5x*2(x^3 + x^2 - x +0)

OpenStudy (mathstudent55):

Answer: 2x^2 5x^5+10x^4 -15x^3+10x^2 = 5x^2(x^3 + 2x^2 ________ What is -15x^3 divided by 5x^2?

OpenStudy (vortish):

5x*2(x^3+5x^2-3x +2)

OpenStudy (mathstudent55):

Answer: -3x 5x^5+10x^4 -15x^3+10x^2 = 5x^2(x^3 + 2x^2 - 3x _______) What is 10x^2 divided by 5x^2? Answer: 2, so 5x^5+10x^4 -15x^3+10x^2 = 5x^2(x^3 + 2x^2 - 3x +2) @vortish Look at you second term inside the parenthses.

OpenStudy (anonymous):

yup thats right haha :)

OpenStudy (vortish):

so i miss typed it it should be 5x^2(x^3+2x^2-3x +2)

OpenStudy (anonymous):

that is correct!

OpenStudy (vortish):

and that is it for the factoring or do we have to simplify it its does not ask

OpenStudy (anonymous):

thats as simplified as it can get.. right @mathstudent55 ?